The mean gasoline computation of 10 cars is 28 liters with a standard deviation of 1.60 liters. Find the point and the interval estimate using 95% confidence level, what is the confidence level



1
Expert's answer
2022-04-26T17:24:13-0400

In statistics, the confidence level indicates the probability, with which the estimation of the location of a statistical parameter (e.g. an arithmetic mean) in a sample survey is also true for the population.

"\ Point\ estimate\,is\ 28"

Z(95%)=1.96 from z value table

"CI=\mu \plusmn z\frac{\sigma}{\sqrt{n}}=28 \plusmn 1.96 \frac{1.6}{\sqrt{10}}=28 \plusmn 0.99"